OpenStudy (anonymous):

For f(x)=1-x/1+x and g(x)=x/1+x and , find the simplified form for f [g(x)] and state the domain. @freckles @radar @ganeshie8

OpenStudy (anonymous):

f[g(x)] means where there is x you put g(x), can you do that

OpenStudy (anonymous):

you put g(x) in x of f(x)

OpenStudy (anonymous):

yes I can, but i am not sure how to simplify the function once i place g(x) into f(g(x)) @Percy*

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[f(x)=\frac{1-x}{1+x}\\[0.4em]\]\[g(x)=\frac{x}{1+x}\]

OpenStudy (anonymous):

\[\large f[g(x)]=\frac{1-\frac{x}{1+x}}{1+\frac{x}{1+x}}\]

OpenStudy (anonymous):

\[\large f[g(x)]=\frac{1-\frac{x}{1+x}}{1+\frac{x}{1+x}}\] \[\large f[g(x)]=\frac{\frac{1+x}{1+x}-\frac{x}{1+x}}{\frac{1+x}{1+x}+\frac{x}{1+x}}\] \[\large f[g(x)]=\frac{\frac{1+x-x}{1+x}}{\frac{1+x+x}{1+x}}\] \[\large f[g(x)]=\frac{1+x-x}{1+x+x}\] \[\large f[g(x)]=\frac{1}{1+2x}\]

OpenStudy (anonymous):

what I don't get is, why is this a calculus question ?

OpenStudy (anonymous):

im inap calculus ab and its part of this long review I just wasn't sure about the simplification its been a really long time since I've seen a composition question

OpenStudy (anonymous):

thanks so much @*-*

OpenStudy (anonymous):

@((*-*))

OpenStudy (anonymous):

I am untaggable, xd

OpenStudy (anonymous):

anyways.... bye, and you welcome!

OpenStudy (anonymous):

I got a notification

OpenStudy (anonymous):

i tried lol thanks

OpenStudy (anonymous):

but, you will never ever ever ever ever ever^(infinity) get my name underlined, and clickable.

OpenStudy (anonymous):

ya i even clicked the link for ur user and it didnt work when I was trying to type it in here

OpenStudy (anonymous):

its ok haha @Nnesha